A certain type of tomato germinates 90% of the time . A backyard farmer planted 25 seeds.
a. What is the probability that exactly 20 seeds germinate.
b. What is the probability that 20 or more seeds germinate
c. What is the mean number of seeds that germinate
d. Determine the variance and standard deviation in the number of seeds that germinate.
1
Expert's answer
2021-07-12T18:28:30-0400

This is a binomial distribution with n=25, p=0.9.

a. "P(X=20)=C_{25}^{20}0.9^{20}(1-0.9)^{25-20}=0.0646."


b. "P(X\ge20)=\sum_{i=20}^{25}C_{25}^i0.9^i(1-0.9)^{25-i}=0.9666."


c. "\mu=np=25*0.9=22.5."


d. "\sigma^2=np(1-p)=25*0.9*0.1=2.25."

"\sigma=\sqrt{2.25}=1.5."